# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v4

@@ Abstract
Forests play a central role in the global carbon cycle  (?). This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery  (?). A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions. Thus, there has been a long-standing effort to link field measurements to global satellite observations  (?).
@@
Recently, with the creation of the BioMassters benchmark dataset  (?) (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice   (?). We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98. In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT  (?) and GSE  (?).
@@
First, a summary of the BioMassters dataset. This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations  (?, ?) informed by extensive field plots. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig. 5b) that would not appear in coarser resolution data. To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
@@
We compared a UNet-based regression model  (?) trained on TESSERA embeddings with SpectralGPT embeddings  (?) combined with a lightweight spatial-temporal encoder  (?) for temporal aggregation and a UPerNet  (?) regression head, because this combination has been found in the Pangaea benchmark  (?) to outperform other foundation models  (?, ?). Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase. Thus, we adopted the Pangaea implementation of their model  (?), which uses preprocessed satellite imagery from the BioMassters dataset. We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline. This model uses a UNet with a temporal attention encoder  (?). Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable. Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
@@
Burned Area Detection
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change  (?, ?, ?). Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends  (?), predicting wildfire occurrence  (?), assessing ecosystem recovery  (?), and establishing effective fire management strategies  (?).
Remote sensing is widely used to detect burned areas and assess severity regionally and globally  (?). Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input  (?, ?). These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals  (?). Most existing methods use optical sensors  (?), which can detect fire-specific spectral signatures but are susceptible to cloud cover  (?), particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens  (?). SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover  (?, ?, ?). Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness  (?, ?). Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products  (?, ?, ?). TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig. 6).
We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6). Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset  (?). UMAP projections  (?) of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters. In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a). To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels. We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites. The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b). Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency. These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps. These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Figure 6C). The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
Stocking Indices in Voluntary Carbon Markets
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere  (?). To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra  (?)) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to assess the performance of the project. Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.  (?). Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see   (?) for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.
A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Figure 7). As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
TESSERA: a Global Open-Source Foundation Model
In this work, we introduce TESSERA, an innovative remote sensing foundation model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data. TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
Figure 3: TESSERA embeddings achieve state-of-the-art performance in crop type classification. The evaluation was conducted on the 2022 Austrian crop dataset. a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio. TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google Satellite Embedding (GSE). b, Patch-based semantic segmentation performance comparison. TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios. c, Weighted F1 score in few-shot learning scenarios. TESSERA maintains a significant performance advantage even with very few training samples per class. Error bars represent the standard deviation over multiple runs. d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation. Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo. Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure. TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies. Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM. All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations. These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration. A 1:1 reference line (blue) is shown in each panel for comparison.
@@
Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California. a, UMAP projection of $20$ TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events. A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance. b, Burned vs. unburned classification performance using random forest classifiers trained on different input representations. TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google Satellite Embeddings (GSE). c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas. As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
@@
Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market (data has been linearly rescaled). Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products. Note that the different products were originally provided in different units and converted to AGB (see   (?) for details).
Acknowledgments
@@
We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
@@
ZF and SK were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge. SJ was funded by a charitable donation from John Bernstein. JB received support from the UKRI CRCRM scheme (grant MR/Y024354). SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology. DAC and SK were funded by the NERC Centre for Landscape Regeneration ( NE/W00495X/1 )
@@
Data curation: ZF, RY, JK, SJ, AM, DAC, SK
@@
Our framework integrates openly available data from the Copernicus Sentinel-2 (S2)  (?) optical and Sentinel-1 (S1)  (?) SAR missions, leveraging the strengths of each sensor. We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22). To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations. For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox  (?).
The Google Satellite Embedding V1 dataset  (?) is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
PRESTO embeddings can be created on demand in Google Earth Engine  (?) using VertexAI  (?).
Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code  (?). To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository  (?), which is the official third-party software repository for Python. This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer  (?) that can be passed directly into downstream tasks. The embeddings themselves are checksummed and indexed using Pooch  (?) to check data integrity and handle retrieval.
The crop type identification downstream task uses the open INVEKOS Austrian crop dataset  (?).
The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset  (?).
The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok  (?).
The AGB downstream task uses the open BioMassters dataset  (?). The code for the winning model of the BioMassters competition is described in  (?) and freely available on the corresponding GitHub page.
The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark  (?).
@@
Supplementary materials
Materials and Methods
Figures S1 to S7
Tables S1- S5
References (97-0)
Supplementary Materials for
TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis
[…]
Equal contribution. Corresponding author. Email: sk818@cam.ac.uk.
This PDF file includes:
Materials and Methods
Figures S1 to S7
Tables S1- S5
Materials and Methods
Model details

### v5

@@ 3.3 Above-ground Forest Biomass Estimation
Forests play a central role in the global carbon cycle [69]. This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [70]. A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions. Thus, there has been a long-standing effort to link field measurements to global satellite observations [71].
@@
Recently, with the creation of the BioMassters benchmark dataset [72] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [73]. We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98. In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [18] and GSE [39].
@@
First, a summary of the BioMassters dataset. This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [74, 75] informed by extensive field plots. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig. 5b) that would not appear in coarser resolution data. To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
@@
We compared a UNet-based regression model [76] trained on TESSERA embeddings with SpectralGPT embeddings [18] combined with a lightweight spatial-temporal encoder [77] for temporal aggregation and a UPerNet [78] regression head, because this combination has been found in the Pangaea benchmark [79] to outperform other foundation models [72, 79]. Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase. Thus, we adopted the Pangaea implementation of their model [79], which uses preprocessed satellite imagery from the BioMassters dataset. We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline. This model uses a UNet with a temporal attention encoder [72]. Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable. Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
@@
3.4 Burned Area Detection
@@
Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California. a, UMAP projection of $20$ TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events. A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance. b, Burned vs. unburned classification performance using random forest classifiers trained on different input representations. TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google AlphaEarth Satellite Embeddings (GSE). c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas. As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [80, 81, 82]. Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [83], predicting wildfire occurrence [84], assessing ecosystem recovery [85], and establishing effective fire management strategies [86].
Remote sensing is widely used to detect burned areas and assess severity regionally and globally [87]. Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [87, 88]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [87]. Most existing methods use optical sensors [87], which can detect fire-specific spectral signatures but are susceptible to cloud cover [89], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [90]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [87, 91, 92]. Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [87, 93]. Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [91, 92, 94]. TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig. 6).
We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Fig. 6). Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [95]. UMAP projections [45] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters. In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Fig. 6a). To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels. We compared three input representations: (1) TESSERA embeddings, (2) Google AlphaEarth Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post- fire Sentinel-2 composites. The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Fig. 6b). Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency. These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps. These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Fig. 6c). The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
3.5 Stocking Indices in Voluntary Carbon Markets
@@
Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market (data has been linearly rescaled). Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products. Note that the different products were originally provided in different units and converted to AGB (see  [38] for details).
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [96]. To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [97]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to assess the performance of the project. Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al. [98]. Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [38] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.
A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Fig. 7). As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
4 Conclusion
In this work, we introduce TESSERA, an innovative remote sensing foundation model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data. TESSERA embeddings establish new performance standards, while our open-source methodology, FAIR-compliant data distribution through version-controlled registries, standard geospatial formats, and the GeoTessera library permit reproducible science and enable use of our work even by non-specialists who lack access to high-performance computing. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
5 Acknowledgments
@@
We gratefully acknowledge support from Mantle Labs and help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
@@
ZF and SK were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge. SJ was funded by a charitable donation from John Bernstein. JB received support from the UKRI CRCRM scheme (grant MR/Y024354). SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology. DAC and SK were funded by the NERC Centre for Landscape Regeneration (NE/W00495X/1). TJ was funded by NERC grant NE/S010750/1. MCL was funded by UKRI and Mantle Labs.
@@
Data collection: TJ
Data analysis: ZF, RY, JK, SJ, AM, TJ, DAC, SK
@@
Our framework integrates openly available data from the Copernicus Sentinel-2 (S2) [99] optical and Sentinel-1 (S1) [100] SAR missions, leveraging the strengths of each sensor. We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22). To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations. For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [101].
The Google Satellite Embedding V1 dataset [39] is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
PRESTO embeddings can be created on demand in Google Earth Engine [102] using VertexAI [103].
Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code [37]. To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository [104], which is the official third-party software repository for Python. This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer [105] that can be passed directly into downstream tasks. The embeddings are indexed using Pooch [106] with per-tile checksums stored in an open Git repository, allowing users to pin specific versions for reproducibility or automatically refresh to the latest embeddings as needed. Additionally, a Parquet index is published daily for rapid spatial queries across the entire dataset.
The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [44].
The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset [95].
The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [66].
The AGB downstream task uses the open BioMassters dataset [72]. The code for the winning model of the BioMassters competition is described in [72] and freely available on the corresponding GitHub page.
The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark [79].
@@
Appendix A Materials and Methods
Table 1: Variable definitions for remote sensing data structure
Variable Dimensions Description
$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
$W,H$ Scalar Spatial dimensions (width, height)
$C$ Scalar Number of spectral channels + polarizations
$T$ Scalar Number of time steps
A.1 Model details